# Democratizing Large Language Models via Personalized Parameter-Efficient Fine-tuning

### v1

@@ 4.1. Experimental Settings
Table 2. Main experiment results on the LaMP benchmark. R-1 and R-L denote ROUGE-1 and ROUGE-L, respectively. $k$ refers to the number of retrieved items, with $k=0$ indicating no retrieval. $\uparrow$ indicates that higher values are better, and $\downarrow$ implies lower values are preferable. For each task, the best score is in bold and the second best is underlined.
@@
Task Metric Non-Personalized RAG PAG RAG+OPPU (Ours) PAG+OPPU (Ours) k=0 Random k=1 k=2 k=4 k=0 k=1 k=0 k=1 k=2 k=4 k=0 k=1 LaMP-1: Personalized Citation Identification Accy $\uparrow$ 0.659 0.650 0.659 0.691 0.691 0.756 0.755 0.683 0.675 0.707 0.723 0.772 0.797 F1 $\uparrow$ 0.657 0.647 0.657 0.689 0.690 0.755 0.755 0.682 0.674 0.705 0.723 0.772 0.794 LaMP-2N: Personalized News Categorization Acc $\uparrow$ 0.787 0.785 0.820 0.832 0.832 0.817 0.817 0.810 0.823 0.834 0.838 0.827 0.831 F1 $\uparrow$ 0.538 0.527 0.598 0.632 0.647 0.623 0.621 0.589 0.615 0.635 0.661 0.648 0.638 LaMP-2M: Personalized Movie Tagging Acc $\uparrow$ 0.478 0.499 0.587 0.598 0.622 0.534 0.587 0.600 0.626 0.634 0.645 0.636 0.648 F1 $\uparrow$ 0.425 0.441 0.512 0.514 0.542 0.476 0.506 0.493 0.531 0.535 0.553 0.536 0.540 LaMP-3: Personalized Product Rating MAE $\downarrow$ 0.223 0.259 0.214 0.214 0.232 0.321 0.223 0.179 0.196 0.214 0.223 0.205 0.143 RMSE $\downarrow$ 0.491 0.590 0.535 0.463 0.535 0.582 0.473 0.443 0.518 0.463 0.526 0.473 0.378 LaMP-4: Personalized News Headline Gen. R-1 $\uparrow$ 0.186 0.187 0.191 0.196 0.198 0.187 0.193 0.904 0.194 0.196 0.199 0.189 0.194 R-L $\uparrow$ 0.167 0.168 0.172 0.176 0.178 0.168 0.173 0.171 0.175 0.177 0.180 0.170 0.175 LaMP-5: Personalized Scholarly Title Gen. R-1 $\uparrow$ 0.476 0.478 0.505 0.510 0.499 0.486 0.516 0.519 0.522 0.511 0.526 0.490 0.525 R-L $\uparrow$ 0.415 0.418 0.445 0.444 0.434 0.429 0.440 0.442 0.457 0.440 0.467 0.428 0.473 LaMP-7: Personalized Tweet Paraphrasing R-1 $\uparrow$ 0.527 0.524 0.568 0.577 0.562 0.542 0.568 0.539 0.579 0.575 0.581 0.542 0.577 R-L $\uparrow$ 0.474 0.474 0.521 0.527 0.514 0.501 0.518 0.483 0.533 0.531 0.528 0.492 0.533
4.1.1. Datasets
@@
We adopt the Large Language Model Personalization (LaMP) benchmark (Salemi et al., 2023) for our experiments, which consists of seven public language model personalization tasks, including four classification tasks (personalized citation identification, personalized movie tagging, personalized producing rating, personalized news categorization) and three generation tasks (personalized news headline generation, personalized scholarly title generation, personalized tweet paraphrasing). [^2] We exclude the LaMP-6: personalized Email subject generation task since it involves private data that we don’t have access to. We mainly focus on the most active uses and select 100 users from the time-based dataset version that have the longest history log as the test set, while all other users are used for base LLM training. Statistics of datasets are in Table 1.
4.1.2. Baselines
We compare our proposed OPPU with the non-personalized baseline and the retrieval-augmented and profile-augmented LLM personalization methods.
  • Retrieval-Augmented Personalization (RAG): We follow the retrieval-augmented personalization method presented in LaMP (Salemi et al., 2023), where the user’s query is augmented with top $k$ retrieved items from the corresponding user’s history corpus. We take $k$=1, 2, 4 in this work.
  • Profile-Augmented Personalization (PAG): This method is taken from Richardson et al. (2023), in which the user’s input sequence would concatenate the user’s profile summarizing the user’s preference and behavior patterns. In our experiments, we generate user profiles using the vicuna-7B (Chiang et al., 2023) model. Moreover, the profile-augmented method could be combined with the retrieval augmentation. In this case, we take the number of retrieval items $k$=1 following the setting of Richardson et al. (2023).
@@ 4.1.2. Baselines
For all baselines, we choose one of the most widely adopted open-source large language model Llama-2-7B (Touvron et al., 2023) as our base LLM and take BM25 (Trotman et al., 2014) for all retrieval operations to ensure efficient and fair comparison.
4.1.3. Evaluation Metrics
@@
Following LaMP (Salemi et al., 2023), we use accuracy and F1-score for classification tasks (LaMP-1: personalized citation identification, LaMP-2N: personalized news categorization, and LaMP-2M: personalized movie tagging), MAE and RMSE for LaMP-3: personalized product rating, and adopt ROUGE-1 and ROUGE-L (Lin, 2004) for text generation tasks (LaMP-4: personalized news headline generation, LaMP-5: personalized scholarly title generation, LaMP-7: personalized tweet paraphrasing). Note that all metrics are the higher the better, except for RMSE and MAE used for the LaMP-3: personalized product rating task.
4.2. Main Results
@@
Table 2 shows the performance on the test set of all seven public tasks in the LaMP benchmark. From the experimental results, we observe that
@@
  • Impact of OPPU. Models equipped with OPPU outperform all corresponding personalization baseline methods across all seven tasks. Notably, in personalized classification tasks, OPPU achieves an average relative improvement of 17.38% and 8.89% on MAE and RMSE metrics on personalized product rating prediction task, as well as 11.87% accuracy and 7.56% F1-score performance gain on personalized movie tagging task. For personalized text generation tasks, we observe a 3.42% and 3.87% enhancement in ROUGE-1 and ROUGE-L scores, respectively, for personalized scholarly title generation.
@@
  • Integrating non-parametric & parametric knowledge. Combining OPPU’s parametric knowledge stored in PEFT parameters and the non-parametric in retrieved items and user profiles, results in notable performance gains. For instance, averaging across all seven tasks, combining retrieval in OPPU will bring 1.93% and 2.48% relative improvement compared with the non-retrieval and non-OPPU yet retrieval version model, respectively. Moreover, integrating OPPU with user profiles would also bring 4.56% and 7.18% performance gain against non-profile and non-OPPU versions, respectively. Overall, integrating non-parametric knowledge from retrieval and user profile and parametric knowledge from personalized PEFT parameter in OPPU generally presents the best performance.
@@
  • Impact of task & history format difference. In tasks like personalized citation identification, there’s a notable discrepancy between the format of user history and the task itself. Here, the user history comprises the user’s publication history, while the task involves a binary classification to identify the correct citation paper. This disparity is similarly observed in the personalized tweet paraphrasing task. In such cases, our proposed OPPU can robustly enhance performance. Specifically, in the personalized citation identification task, OPPU contributes to a notable increase of 3.48% in accuracy and 3.52% in the F1-score. This improvement is attributed to the provision of personalized context knowledge in a parametric manner via private PEFT.
@@
  • Number of retrieved items. Our experimental results generally indicate that an increase in the number of retrieved items correlates with improved performance. However, we also observe that some data points don’t fit this trend, and we hypothesize that this inconsistency may arise from the retrieved items introducing noise and irrelevant behavior patterns, potentially complicating the model’s process of understanding user preferences.
@@
Table 3. Performance under user behavior shift, where we remove the user behavior history highly similar to the query at hand. $k$ denotes the number of retrieved history items, and $k=0$ means non-retrieval. Armed with irrelevant user history, the retrieval-only method falls short and performs close to the non-personalized baseline, while OPPU shows stronger generalizability in the user behavior shift scenario.
@@ 7 R-1 R-L R-1 R-L R-1 R-L R-1 R-L
4.3. Performance under User Behavior Shift
@@
Recent studies have shown that retrieval-augmented generation methods tend to underperform when the retrieved corpus does not contain highly relevant documents (Petroni et al., 2020; Shi et al., 2023; Gupta et al., 2024). This issue often emerges in personalization contexts where the retrieval corpus is confined to a specific user’s behavior history and the user does not have a history of behavior closely matching their current queries. In our experiment, we simulate such a scenario where there is little or no overlap between the user’s behavior history corpus and their current query. Specifically, we employ the encoder-only language model DeBERTa-v3 (He et al., 2022) to extract features for the user’s historical behaviors and the query, then compute the cosine similarity between the query and all historical items to assess their relevance. By ranking the historical behaviors, we select the top 100 items with the lowest relevance scores as irrelevant user history.
@@
As is demonstrated in Table 3, limiting user history to less relevant results in a marked decline in the performance of retrieval-based methods, often close to the performance of non-personalized approaches. Our proposed OPPU shows stronger robustness and generalization to these less relevant history behaviors and would even outperform using all user history items for private PEFT training. Moreover, the integration of both parametric and non-parametric knowledge (OPPU, k=1) exhibits enhanced robustness in personalized text generation tasks. In contrast, models utilizing only parametric user knowledge (OPPU, k=0) perform better in personalized text classification tasks.
@@
Figure 3. Model performance on personalized movie tagging and personalized tweet paraphrasing for users with different numbers of behavior history.
Figure 4. Cosine similarities between personal PEFT parameters under personalized text classification and generation tasks.
4.4. Modeling Users with Different Active Levels
@@
In the main experiment, we concentrate on modeling highly active users. However, it is worth noting that a significant number of users exhibit lower levels of activity, resulting in a comparatively short behavior history. To investigate the impact of user activity levels quantified by the number of historical behavioral items on model performance, we randomly chose 20 users from each range of active levels. As illustrated in Figure 3, equipped with OPPU, LLMs generally outperform the baseline methods across different user activity levels. Specifically, 1) the longer the user history length, the superiority of retrieval+OPPU over the baseline is generally larger. 2) The inclusion of non-parametric user knowledge via retrieval results in performance improvements compared to methods without retrieval. 3) Integrating the parametric knowledge in OPPU and non-parametric knowledge in retrieval generally shows the strongest performance over users across different active levels.
@@
Figure 5. Performance of OPPU and retrieved-only baseline when the number of retrieved items $k$ increases.
4.5. Impact of Retrieved History Items
@@
In this study, we alter the number of retrieved items of both retrieval-only baseline and retrieval+OPPU to gain a better understanding of the integration of non-parametric and parametric user knowledge. Figure 5 illustrates that as we increase the number of retrieved historical behavior items, both the retrieval-only baselines and the retrieval+OPPU approaches show improved performance. Interestingly, we observe that as the number of retrieved items $k$ becomes larger, the performance difference between the retrieval-only baseline and retrieval+OPPU narrows. This trend could be attributed to the longer logs of user behavior history in non-parametric prompts, which reduce the gap between the comprehensive user behavior history encapsulated in personalized PEFT parameters and the non-parametric user knowledge included in the prompts.
4.6. Similarities Between Personalized PEFTs
To gain a better understanding of how users’ behavior biases are encapsulated within their private PEFT parameters, we analyze the cosine similarities between these parameters across different users, as illustrated in Figure 4. Specifically, we select two representative tasks from text classification and generation categories respectively, then we compute the cosine similarities on the 100 users’ PEFT parameters in the test set. As shown in Figure 4, we observe that the private PEFT similarities generally range from 0.4 to 0.7. Interestingly, the personalized scholarly title generation task exhibits the highest average similarities, likely due to task-specific characteristics that entail less personal bias. Besides the absolute values of these similarities, the relative differences among various users provide additional insights. In personalized text classification tasks, the similarities tend to exhibit more variance, suggesting that some users have higher similarities compared to others. In contrast, the similarities in personalized text generation tasks remain relatively uniform. This pattern leads us to speculate that personal preferences in text generation tasks are more challenging to categorize, making it harder to distinctly group users based on their preferences. On the other hand, preferences in text classification tasks appear to be more identifiable and classifiable.
@@ 4.6. Similarities Between Personalized PEFTs
Table 4. Performance of OPPU with different ablated versions of user history configurations. $k$ refers to the number of retrieved items, and $k=0$ denotes non-retrieval. The best score is in bold and the second best is underlined.
@@
$\checkmark$ 0.530 0.488 0.486 0.437 0.624 0.539
$\checkmark$ 0.567 0.514 0.499 0.44 0.634 0.548
$\checkmark$ $\checkmark$ 0.587 0.512 0.600 0.493 0.626 0.531
@@
$\checkmark$ 0.493 0.422 0.497 0.434 0.495 0.449
$\checkmark$ 0.475 0.425 0.489 0.430 0.492 0.429
$\checkmark$ $\checkmark$ 0.505 0.445 0.519 0.442 0.522 0.457
@@
Figure 6. We present a case study on a specific query from a user and analyze the responses generated by a non-personalized model in the personalized movie tagging task, a retrieval-augmented personalization model, and our proposed OPPU. It is shown that the retrieval-augmented personalization method can be easily distracted by less relevant user behavior history. In contrast, our OPPU demonstrates a more effective and comprehensive ability to capture the user’s behavior patterns.
@@
Figure 7. Performance of OPPU on personalized movie tagging and personalized scholarly title generation tasks when equipped with different PEFT methods. We find that a larger proportion of trainable parameters generally results in better personalization performance.
4.7. Robustness against Task Formats
@@
Our main results demonstrate that even with a history corpus that does not strictly follow the task format, our plug-in OPPU would bring significant performance improvement. In this study, we ablate the history format to test the robustness against user history format on personalized movie tagging (LaMP-2M) and personalized scholarly title generation (LaMP-5), from text classification and generation category respectively. Specifically, in both tasks, each user history item consists of both input and output aligned with the user query $x_{u}$ and output $y_{u}$ at hand. We ablate the history behavior items from the input and output side respectively and compare them with the retrieval baseline to test OPPU’s robustness against the history in mismatched format.
@@
As is shown in Table 4, even with incomplete user behavior history that is not aligned with task format, OPPU still achieves relatively close performance with full history in text generation task. In the news categorization, LLM struggles to correctly classify with only parametric knowledge. However, integrating with retrieval augmentation, OPPU shows strong and robust performance, which can even outperform the full model tuned on the complete pairs of user history data. Overall, experimental results reveal that the integration of both non-parametric and parametric knowledge can make a robust model with different formats of user history.
4.8. On PEFT Method Choices

### v2

@@ 2.2 Base LLMs Task Adaption
where $\mathrm{CE}$ denotes the cross entropy loss function, $\phi_{t}$, $\phi_{r}$, and $\phi_{p}$ denote prompt construction function for base, retrieval-augmented, and profile-augmented LLM. The retrieved user history $\mathcal{D}_{u}=\mathcal{R}(q_{u},\mathcal{H}_{u},k)$ denotes the top-$k$ user history from retriever $\mathcal{R}$. $s_{u}=\mathrm{LLM}(\mathcal{H}_{u})$ is a textual user profile generated by an instruction-tuned LLM, e.g., Vicuna Chiang et al. (2023), based on user history.
To make this process more computationally efficient, we adopt the low-rank adaptation (LoRA) Hu et al. (2021) for base LLM task adaption that only updates about 0.5% external parameters compared to the total LLM parameter size. After training, LoRA parameters are merged into the base model, equipping LLMs with task capabilities.
3 One PEFT Per User (OPPU)
Once the base model for task adaption is obtained, users can only access the base model parameters and their personal behavior history data, controlling privacy risks. This section introduces personalized LLMs for target users through parametric PEFT and integrates non-parametric knowledge such as retrieval and profile augmentation. For each user, we plug a personal trainable PEFT module (LoRA by default) $\Delta\Theta_{u}^{\mathrm{(B)}}$, $\Delta\Theta_{u}^{\mathrm{(R)}}$, $\Delta\Theta_{u}^{\mathrm{(P)}}$ to corresponding base LLM under three settings to obtain personalized LLM $\Theta_{u}^{\mathrm{(B)}}$, $\Theta_{u}^{\mathrm{(R)}}$, and $\Theta_{u}^{\mathrm{(P)}}$, while base LLM parameters $\Theta^{\mathrm{(B)}}$, $\Theta^{\mathrm{(R)}}$, $\Theta^{\mathrm{(P)}}$ are frozen.
@@ 3 One PEFT Per User (OPPU)
We then use the user data $\mathcal{H}_{u}$ for LLM fine-tuning to learn the personalized PEFT parameters. The training objectives for user $u$ under base, retrieval-augmented, and profile-augmented settings are:
[…]
where $\mathcal{D}_{u}^{<t(x_{u})}=\mathcal{R}(\phi_{t}(x_{u}),\mathcal{H}_{u}^{<t(x_{u})},k)$, $\mathcal{H}_{u}^{<t(x_{u})}$ is restricted to user $u$’s past behavior history that occurred before $x_{u}$.
User behavior history often does not align neatly with the query format. For example, in personalized tweet paraphrasing tasks, where the input is a text sequence $q_{u}$ and the output is the paraphrased tweet $r_{u}$, the history $\mathcal{H}_{u}$ only includes historical tweets. In scenarios where user history does not directly aligned with the specific task format, denoted as $\mathcal{H}_{u}=\{x_{u}\}$, we replace the user history output $y_{u}$ in personal PEFT training objectives $\mathcal{L}_{u}^{\mathrm{(B)}}$, $\mathcal{L}_{u}^{\mathrm{(R)}}$, $\mathcal{L}_{u}^{\mathrm{(P)}}$ with right-shifted history $x_{u}^{\prime}$ for unsupervised next token prediction.
By optimizing personal PEFT parameters with the objectives mentioned above, OPPU comprehensively capture the user behavior patterns in PEFT parameters $\Delta\Theta_{u}^{\mathrm{(B)}}$, $\Delta\Theta_{u}^{\mathrm{(R)}}$, $\Delta\Theta_{u}^{\mathrm{(P)}}$, creating personalized LLMs owned by users. We envision the proposed OPPU as a versatile LLM personalization framework, where each user possesses their own PEFT parameters that contain personal behavior history and preferences. By plugging their personal PEFT parameters into the base LLMs, users can get their personalized LLMs, while achieving a better understanding and generalization of users’ preferences from the parametric dimension.
4 Experimental Settings
@@
Table 1: Main experiment results on the LaMP benchmark. R-1 and R-L denote ROUGE-1 and ROUGE-L, respectively. $k$ refers to the number of retrieved items, with $k=0$ indicating no retrieval. $\uparrow$ indicates that higher values are better, and $\downarrow$ implies lower values are preferable. For each task, the best score is in bold and the second best is underlined.‘’ indicates significant improvement against counterparts without OPPU.
@@
Task Metric Non-Personalized RAG PAG RAG+OPPU (Ours) PAG+OPPU (Ours) k=0 Random k=1 k=2 k=4 k=0 k=1 k=0 k=1 k=2 k=4 k=0 k=1 LaMP-1: Personalized Citation Identification Acc $\uparrow$ .659 .650 .659 .691 .691 .756 .755 .683 .675 .707 .723 .772 .797 F1 $\uparrow$ .657 .647 .657 .689 .690 .755 .755 .682 .674 .705 .723 .772 .794 LaMP-2N: Personalized News Categorization Acc $\uparrow$ .787 .785 .820 .832 .832 .817 .817 .810 .823 .834 .838 .827 .831 F1 $\uparrow$ .538 .527 .598 .632 .647 .623 .621 .589 .615 .635 .661 .648 .638 LaMP-2M: Personalized Movie Tagging Acc $\uparrow$ .478 .499 .587 .598 .622 .534 .587 .600 .626 .634 .645 .636 .648 F1 $\uparrow$ .425 .441 .512 .514 .542 .476 .506 .493 .531 .535 .553 .536 .540 LaMP-3: Personalized Product Rating MAE $\downarrow$ .223 .259 .214 .214 .232 .321 .223 .179 .196 .214 .223 .205 .143 RMSE $\downarrow$ .491 .590 .535 .463 .535 .582 .473 .443 .518 .463 .526 .473 .378 LaMP-4: Personalized News Headline Gen. R-1 $\uparrow$ .186 .187 .191 .196 .198 .187 .193 .191 .194 .196 .199 .189 .194 R-L $\uparrow$ .167 .168 .172 .176 .178 .168 .173 .171 .175 .177 .180 .170 .175 LaMP-5: Personalized Scholarly Title Gen. R-1 $\uparrow$ .476 .478 .505 .510 .499 .486 .516 .519 .522 .511 .526 .490 .525 R-L $\uparrow$ .415 .418 .445 .444 .434 .429 .440 .442 .457 .440 .467 .428 .473 LaMP-7: Personalized Tweet Paraphrasing R-1 $\uparrow$ .527 .524 .568 .577 .562 .542 .568 .539 .579 .575 .581 .542 .577 R-L $\uparrow$ .474 .474 .521 .527 .514 .501 .518 .483 .533 .531 .528 .492 .533
Table 2: Dataset statistics: We report average sequence length in terms of number of tokens. #Q is the number of queries, Lin and Lout are the average length of input and output sequence respectively, and #History is the number of adopted items. To save space, task names can be found in Table 1.
Task in LaMP Base LLM Training Personal PEFT Training #Q Lin Lout #Q #History Lin Lout 1 7,919 51.3 1.0 123 317.5 52.0 1.0 2M 3,181 92.1 1.4 3,302 55.6 92.6 2.0 2N 3,662 68.2 1.3 6,033 219.9 63.5 1.1 3 22,388 128.7 1.0 112 959.8 211.9 1.0 4 7,275 33.9 9.2 6,275 270.1 25.2 11.1 5 16,075 162.1 9.7 107 442.9 171.6 10.3 7 14,826 29.7 18.3 109 121.2 29.4 18.0
Datasets
@@
We use data from the Large Language Model Personalization (LaMP) benchmark Salemi et al. (2023), which includes seven public language model personalization tasks: four classification tasks and three generation tasks. [^2] We exclude LaMP-6 as it involves private data that we cannot access. To promote LLM ownership, we emphasize the need for users to contribute extensive historical data for personalizing their model. Therefore, we focus on the most active users, selecting 100 users with the longest history logs from the time-based dataset version as the test set, while using all other users for base LLM training. Dataset statistics are presented in Table 2.
Baselines
@@
We compare our proposed OPPU with the non-personalized baseline and the retrieval-augmented (RAG) and profile-augmented (PAG) LLM personalization methods. For all baselines and OPPU, we choose one of the most widely adopted open-source LLM Llama-2-7B Touvron et al. (2023) as our base LLM and take BM25 Trotman et al. (2014) for all retrieval operations to ensure efficient and fair comparison. [^3] Baselines and hyperparameter details are presented in Appendix C and A to facilitate further research.
Evaluation Metrics
@@
Following LaMP Salemi et al. (2023), we use accuracy and F1-score for classification tasks (LaMP-1, LaMP-2N, and LaMP-2M), MAE and RMSE for LaMP-3, and adopt ROUGE-1 and ROUGE-L Lin (2004) for text generation tasks (LaMP-4, LaMP-5, LaMP-7). Note that all metrics are the higher the better, except for RMSE and MAE used for the LaMP-3.
5 Results
@@
Table 1 shows the performance on the test set for all seven public tasks in the LaMP benchmark, we have observations as follows.
OPPU brings universal improvement.
@@
Models equipped with OPPU outperform all baseline personalization methods across all seven tasks. Notably, in personalized classification tasks, OPPU achieves an average relative improvement of 17.38% in MAE and 8.89% in RMSE for personalized product rating prediction. Additionally, it shows an 11.87% improvement in accuracy and 7.56% in F1-score for personalized movie tagging. For personalized text generation tasks, OPPU enhances ROUGE-1 and ROUGE-L scores by 3.42% and 3.87%, respectively, in personalized scholarly title generation.
Integrating non-parametric and parametric knowledge performs the best.
@@
Combining OPPU’s parametric knowledge stored in PEFT parameters and the non-parametric in retrieved items and user profiles, results in notable performance gains. For instance, averaging across all seven tasks, combining retrieval in OPPU will bring 1.93% and 2.48% relative improvement compared with the non-retrieval and non-OPPU yet retrieval version model, respectively. Moreover, integrating OPPU with user profiles would also bring 4.56% and 7.18% performance gain against non-profile and non-OPPU versions, respectively. Overall, combining non-parametric retrieval and profile knowledge with parametric PEFT knowledge in OPPU delivers the best performance.
Performance w.r.t. difference between task and history format.
@@
In tasks like personalized citation identification, there is a notable discrepancy between the user history format and the task itself. Here, the user history comprises the user’s publication history, while the task involves binary classification to identify the correct citation paper. This disparity is also seen in the personalized tweet paraphrasing task. In these cases, OPPU significantly enhances performance. Specifically, for personalized citation identification, OPPU increases accuracy by 3.48% and F1-score by 3.52%, thanks to personalized context knowledge provided through personal PEFT.
The more retrieved items, the better performance.
@@
Our experimental results generally indicate that an increase in the number of retrieved items correlates with improved performance. However, we also observe that some data points don’t fit this trend, and we hypothesize that this inconsistency may arise from the retrieved items introducing noise and irrelevant behavior patterns, potentially complicating the model’s process of understanding user preferences.
@@
Table 3: Performance under user behavior shift, where we remove the user behavior history highly similar to the query at hand. $k$ denotes the number of retrieved history items, and $k=0$ means non-retrieval. Armed with irrelevant user history, the retrieval-only method falls short and performs close to the non-personalized baseline, while OPPU shows stronger generalizability in the user behavior shift scenario.
@@ 7 R-1 R-L R-1 R-L R-1 R-L R-1 R-L
6 Analysis
Performance under User Behavior Shift
@@ 6 Analysis
Recent studies have shown that retrieval-augmented generation methods tend to underperform when the retrieved corpus does not contain highly relevant documents Shi et al. (2023); Gupta et al. (2024). This problem is common in personalization contexts where the user’s behavior history does not closely match their current queries. To simulate this scenario, we use DeBERTa-v3 He et al. (2022) to extract features from the user’s historical behaviors and current query, computing cosine similarity to assess relevance. We then rank the historical behaviors and select the top 100 items with the lowest relevance scores as irrelevant user history.
@@
Table 3 shows that limiting user history to less relevant items significantly reduces the performance of retrieval-based methods, often aligning with non-personalized approaches. In contrast, OPPU demonstrates stronger robustness and generalization to less relevant history, even outperforming models trained with all user history items. Additionally, the combination of parametric and non-parametric knowledge (OPPU, $k$=1) enhances robustness in personalized text generation tasks, while models using only parametric knowledge (OPPU, $k$=0) perform better in personalized text classification tasks.
@@
Figure 3: Model performance on personalized movie tagging and personalized tweet paraphrasing for users with different numbers of behavior history.
Modeling Users with Different Active Levels
@@
In our main experiment, we focus on highly active users. However, many users exhibit lower activity levels, resulting in shorter behavior histories. To examine the impact of user activity levels on model performance, we randomly selected 20 users from each activity range. Figure 3 shows that LLMs equipped with OPPU consistently outperform baseline methods across various activity levels. Key observations include: 1) The longer the user history, the more pronounced the superiority of retrieval + OPPU over baselines. 2) Including non-parametric user knowledge via retrieval improves performance compared to methods without retrieval. 3) Integrating parametric knowledge in OPPU with non-parametric knowledge from retrieval yields the strongest performance across different user activity levels.
@@
Figure 4: Performance of OPPU and retrieved-only baseline when the number of retrieved items $k$ increases.
Performance w.r.t. Retrieved History Items k
@@
In this study, we alter the number of retrieved items of both retrieval-only baseline and retrieval+OPPU to gain a better understanding of the integration of non-parametric and parametric user knowledge. Figure 4 illustrates that as we increase the number of retrieved historical behavior items, both the retrieval-only baselines and the retrieval+OPPU approaches show improved performance. Interestingly, we observe that as the number of retrieved items $k$ becomes larger, the performance difference between the retrieval-only and retrieval+OPPU narrows. This trend could be attributed to the longer logs of user behavior history in non-parametric prompts, which reduce the gap between the comprehensive user behavior history encapsulated in personalized PEFT parameters and the non-parametric user knowledge included in the prompts.
@@
Table 4: Performance of OPPU with different ablated versions of user history configurations. $k$ refers to the number of retrieved items, and $k=0$ denotes non-retrieval. The best score is in bold and the second best is underlined.
@@
$\checkmark$ .530 .488 .486 .437 .624 .539
$\checkmark$ .567 .514 .499 .440 .634 .548
$\checkmark$ $\checkmark$ .587 .512 .600 .493 .626 .531
@@
$\checkmark$ .493 .422 .497 .434 .495 .449
$\checkmark$ .475 .425 .489 .430 .492 .429
$\checkmark$ $\checkmark$ .505 .445 .519 .442 .522 .457
@@
Figure 5: Case study in the personalized movie tagging task. It is shown that the retrieval-augmented personalization method can be easily distracted by less relevant user behavior history. In contrast, our OPPU demonstrates a more effective and comprehensive ability to capture the user’s behavior patterns.
@@
Figure 6: Performance of OPPU on personalized movie tagging and personalized scholarly title generation tasks when equipped with different PEFT methods. We find that a larger proportion of trainable parameters generally results in better personalization performance.
Robustness against Task Formats
@@
Our main results demonstrate that OPPU significantly improves performance even when the user history corpus does not strictly follow the task format. We tested this robustness by ablating the history format in personalized movie tagging (LaMP-2M) and personalized scholarly title generation (LaMP-5) tasks, covering both text classification and generation categories. In both tasks, each user history item consists of input and output aligned with the user query $x_{u}$ and output $y_{u}$. We ablated history behavior items from the input and output sides, comparing them with the retrieval baseline to test OPPU’s robustness against mismatched formats.
@@
Shown in Table 4, OPPU achieves performance close to that with full history in the text generation task, even with incomplete user behavior history. In news categorization, LLMs struggle with only parametric knowledge, but integrating retrieval augmentation, OPPU shows robust performance, outperforming models tuned on complete user history data. Overall, results reveal that combining non-parametric and parametric knowledge makes OPPU robust to different user history formats.
On PEFT Method Choices